Two electrons are passing 21 mm apart. What is the electric repulsive force that they exert on each other?
Select one:
a. 2.0 N
b. 2.0
Answer #1

F=kq2/r2=9*109*(1.6*10-19)2/(21*10-3)2=5.2*10-25N

c) is correct